# I/O in Machine Learning Applications on HPC Systems: A 360-degree Survey

### v1

@@ Acknowledgements.
- Smilkov et al. (2019) Daniel Smilkov, Nikhil Thorat, Yannick Assogba, Ann Yuan, Nick Kreeger, Ping Yu, Kangyi Zhang, Shanqing Cai, Eric Nielsen, David Soergel, Stan Bileschi, Michael Terry, Charles Nicholson, Sandeep N. Gupta, Sarah Sirajuddin, D. Sculley, Rajat Monga, Greg Corrado, Fernanda B. Viégas, and Martin Wattenberg. 2019. TensorFlow.js: Machine Learning for the Web and Beyond. arXiv:1901.05350 [cs.LG]
- Smith et al. (2020) Samuel Smith, Erich Elsen, and Soham De. 2020. On the Generalization Benefit of Noise in Stochastic Gradient Descent. In Proceedings of the 37th International Conference on Machine Learning (Proceedings of Machine Learning Research, Vol. 119), Hal Daumé III and Aarti Singh (Eds.). PMLR, Cornell University, 9058–9067. https://proceedings.mlr.press/v119/smith20a.html
- Snyder et al. (2016a) Shane Snyder, Philip Carns, Kevin Harms, Robert Ross, Glenn K. Lockwood, and Nicholas J. Wright. 2016a. Modular HPC I/O Characterization with Darshan. In 2016 5th Workshop on Extreme-Scale Programming Tools (ESPT). 2016 5th workshop on extreme-scale programming tools (ESPT), Salt Lake City, UT, USA, 9–17. https://doi.org/10.1109/ESPT.2016.006
- Snyder et al. (2016b) Shane Snyder, Philip Carns, Kevin Harms, Robert Ross, Glenn K. Lockwood, and Nicholas J. Wright. 2016b. Modular HPC I/O Characterization with Darshan. In Proceedings of the 5th Workshop on Extreme-Scale Programming Tools (Salt Lake City, Utah) (ESPT ’16). IEEE Press, 9–17.
@@
- Sun et al. (2022) Xiaoyang Sun, Wei Wang, Shenghao Qiu, Renyu Yang, Songfang Huang, Jie Xu, and Zheng Wang. 2022. STRONGHOLD: Fast and Affordable Billion-Scale Deep Learning Model Training. In SC22: International Conference for High Performance Computing, Networking, Storage and Analysis. 1–17. https://doi.org/10.1109/SC41404.2022.00076
@@
- Villalobos et al. (2022) Pablo Villalobos, Jaime Sevilla, Tamay Besiroglu, Lennart Heim, Anson Ho, and Marius Hobbhahn. 2022. Machine Learning Model Sizes and the Parameter Gap. arXiv:2207.02852 [cs.LG]
- Wan et al. (2020) Lipeng Wan, Matthew Wolf, Feiyi Wang, Jong Youl Choi, George Ostrouchov, Jieyang Chen, Norbert Podhorszki, Jeremy Logan, Kshitij Mehta, Scott Klasky, et al. 2020. I/O Performance Characterization and Prediction through Machine Learning on HPC Systems. In CUG2020. CUG, Dallas Convention Center Room A309-A310 650 S Griffin St Dallas, TX 75202, 10.
@@
- xie et al. (2022) jian xie, Jingwei Xu, Guochang Wang, Yuan Yao, Zenan Li, Chun Cao, and Hanghang Tong. 2022. A Deep Learning Dataloader with Shared Data Preparation. In Advances in Neural Information Processing Systems, S. Koyejo, S. Mohamed, A. Agarwal, D. Belgrave, K. Cho, and A. Oh (Eds.), Vol. 35. Curran Associates, Inc., Virtual, 17146–17156. https://proceedings.neurips.cc/paper_files/paper/2022/file/6d538a6e667960b168d3d947eb6207a6-Paper-Conference.pdf
@@
- Zhu et al. (2018) Maohua Zhu, Youwei Zhuo, Chao Wang, Wenguang Chen, and Yuan Xie. 2018. Performance Evaluation and Optimization of HBM-Enabled GPU for Data-Intensive Applications. IEEE Transactions on Very Large Scale Integration (VLSI) Systems 26, 5 (2018), 831–840. https://doi.org/10.1109/TVLSI.2018.2791442

### v2

@@ Acknowledgements.
- Snyder et al. (2016) Shane Snyder, Philip Carns, Kevin Harms, Robert Ross, Glenn K. Lockwood, and Nicholas J. Wright. 2016. Modular HPC I/O Characterization with Darshan. In Proceedings of the 5th Workshop on Extreme-Scale Programming Tools (Salt Lake City, Utah) (ESPT ’16). IEEE Press, 9–17.
- Song et al. (2021) Xinying Song, Alex Salcianu, Yang Song, Dave Dopson, and Denny Zhou. 2021. Fast WordPiece Tokenization. arXiv:2012.15524 [cs.CL] https://arxiv.org/abs/2012.15524
@@
- Thakur (2024) Rajeev Thakur. 2024. AuroraGPT: A Large-Scale Foundation Model for Advancing Science. https://doi.org/10.5281/zenodo.13345059
@@
- Xu et al. (2021) Cong Xu, Suparna Bhattacharya, Martin Foltin, Suren Byna, and Paolo Faraboschi. 2021. Data-Aware Storage Tiering for Deep Learning. In 2021 IEEE/ACM Sixth International Parallel Data Systems Workshop (PDSW). 23–28. https://doi.org/10.1109/PDSW54622.2021.00009
- Xu et al. (2023) Chenhao Xu, Youyang Qu, Yong Xiang, and Longxiang Gao. 2023. Asynchronous federated learning on heterogeneous devices: A survey. Computer Science Review 50 (2023), 100595. https://doi.org/10.1016/j.cosrev.2023.100595
@@
- Zhu et al. (2015) Yukun Zhu et al. 2015. Aligning Books and Movies: Towards Story-Like Visual Explanations by Watching Movies and Reading Books. In 2015 IEEE International Conference on Computer Vision (ICCV). 19–27. https://doi.org/10.1109/ICCV.2015.11